# AdvMT: Adversarial Motion Transformer for Long-term Human Motion Prediction

### 10 Jan 2024 v1

@@ Abstract
To achieve seamless collaboration between robots and humans in a shared environment, accurately predicting future human movements is essential. Human motion prediction has traditionally been approached as a sequence prediction problem, leveraging historical human motion data to estimate future poses. Beginning with vanilla recurrent networks, the research community has investigated a variety of methods for learning human motion dynamics, encompassing graph-based and generative approaches. Despite these efforts, achieving accurate long-term predictions continues to be a significant challenge. In this regard, we present the Adversarial Motion Transformer (AdvMT), a novel model that integrates a transformer-based motion encoder and a temporal continuity discriminator. This combination effectively captures spatial and temporal dependencies simultaneously within frames. With adversarial training, our method effectively reduces the unwanted artifacts in predictions, thereby ensuring the learning of more realistic and fluid human motions. The evaluation results indicate that AdvMT greatly enhances the accuracy of long-term predictions while also delivering robust short-term predictions.
@@ I Introduction
The ability to accurately predict human motion is a crucial aspect in advancing human-robot interactions, an area that has garnered significant attention in both academic and industrial circles. At its core, the problem of human motion prediction aims to understand and forecast the complex and dynamic movements of humans, which is essential for enabling robots to interact safely and effectively with their human counterparts. This technology has broad applications, ranging from assistive robotics in healthcare, where robots aid in patient care, to advanced manufacturing settings, where collaborative robots work alongside human workers. However, the task of predicting human motion is challenging due to a multitude of influencing factors, including individual physical differences, psychological states, and the surrounding environment. These factors collectively complicate the development of realistic and reliable human motion prediction systems, a critical step for ensuring the smooth and harmonious integration of robots in a human-centric environment.
@@ II-A Long-term human motion prediction
Human motion prediction has been a cornerstone in the field of computer vision and robotics, witnessing significant evolution over the years. The initial phase of research in this area primarily utilized RNNs with encoder-decoder models [1, 2, 3], but these models faced challenges in handling complex human motion dependencies. This led to a shift towards CNNs, which provided improved extraction of spatial joint connection information [5, 14, 6]. Later, the exploration of GCNs brought enhanced capabilities in anatomical relationship modeling [7, 8, 15]. However, much of the research has concentrated mainly on short-term predictions, leaving a notable gap in long-term prediction accuracy.
@@ III Adversarial Motion Transformer (AdvMT)
The overview of the Adversarial Motion Transformer (AdvMT), a system comprised of two main branches, is illustrated in Fig. 2. The motion encoder branch interprets the input motion history and encodes the local and global human joint dependencies. The discriminator branch complements this by refining the predictions from the motion encoder branch to ensure the generation of realistic and consistent human motion.
@@ III-D Loss function
TABLE III: Long-term prediction results for H3.6M dataset.
Walking Smoking Purchases Sitting
Time (ms) 1.5k 2k 1.5k 2k 1.5k 2k 1.5k 2k
HisRepeat [8] 64.6 73.4 87.7 101.8 159.1 172.2 156.6 178.9
AdvMT (ours) 62.4 73.1 91.4 99.0 140.4 153.1 151.0 164.2
@@ IV Experiments
Dataset: In our experiments, we conducted our model evaluation using the Human3.6M dataset, widely recognized as a benchmark in the field of human motion prediction. This extensive database contains over 3.6 million 3D poses, recorded with 7 actors performing 15 different types of actions. For training and evaluation, we downsampled the motion data to 25 frames per second. In alignment with the protocol outlined in [2], we used subjects S1, S6, S7, S8, S9, and S11 for training, and S5 is designated for testing. Similar to [8], we report our future motion prediction results on 256 sub-sequences per action.
Comparison with other methods: In our study, we focused on training our method for 3D joint position prediction. We compared its performance with other state-of-the-art methods trained for the same motion representation, as detailed in Tables I and II. The primary metric used for evaluation is the MPJPE in millimeters, aligning with the standards used by other methods [8, 18]. We trained the model using an input sequence of 2 seconds (50 frames) to generate predictions for the next 1 second (25 frames). However, our motion encoder branch was trained auto-regressively, enabling it to predict motions extending beyond 1 second (see Table III and Fig. 4). The results for BiTGAN [18] are acquired from the published paper, while the HisRepeat [8] method is evaluated with the provided code and the trained model.
@@ V-A Architecture
An ablation study was conducted to assess the effectiveness of our proposed AdvMT architecture. We compared the performance of a full Transformer network, comprising both encoder and decoder layers with separate self-attention and multi-head attention mechanisms, against our modified architecture. Originally intended for sequence-to-sequence tasks, the full Transformer architecture was found less effective for human motion prediction compared to our adaptation, which solely utilizes the encoder layer. This suggests that the decoder layer might add unnecessary complexity, hindering the ability to capture human motion dynamics accurately.
Further ablation studies highlighted the critical role of the discriminator branch in enhancing the realism and temporal consistency of generated motion. This branch serves as a feedback mechanism for the motion encoder, guiding it to correct and refine its predictions. Without it, the motion encoder tended to yield less realistic and plausible motions. The discriminator branch functions as a critic, ensuring the predicted motions are not only realistic but also align with real-world motion patterns. Its absence leads to the generation of unrealistic or inconsistent motion sequences, particularly in complex human motions where accuracy in joint angles and movements is crucial. By integrating the adversarial training, our model is compelled to generate more lifelike and consistent motion sequences.
@@ V-B Loss function
Regarding the loss function, we found that using only the bone length error as a loss term with the vanilla MPJPE loss led the model to predict zero-velocity motion for long-term prediction tasks. As the model learns to minimize only the bone length error, leading it to generate static poses. To mitigate this, incorporating the discriminator loss term helped to penalize unrealistic human motion predictions and encouraged the motion encoder to generate more realistic and plausible human motion. We conducted an ablation study to investigate the effectiveness of our modified loss function, which includes a discriminator loss and a bone error loss in addition to the regular MPJPE loss. We compared the performance of our model with the modified loss function against a baseline model that only used the regular MPJPE loss (see Table IV).
The results show that the combination of all three losses achieved the best performance while using only one of the losses resulted in lower performance. The bone error loss contributed the most to the overall performance improvement, followed by the discriminator loss. Our modified loss function, which includes the MPJPE loss, bone length error, and a temporal continuity discriminator loss, is effective in improving the quality of human motion prediction.
TABLE IV: The ablation study results on H3.6M dataset.

### 19 Feb 2024 v2

@@ Abstract
Human motion prediction has traditionally been approached as a sequence prediction problem, leveraging historical human motion data to estimate future poses. Beginning with vanilla recurrent networks, the research community has investigated a variety of methods for learning human motion dynamics, encompassing graph-based and generative approaches. Despite these efforts, achieving accurate long-term predictions continues to be a significant challenge. In this regard, we present the Adversarial Motion Transformer (AdvMT), a novel model that integrates a transformer-based motion encoder and a temporal continuity discriminator. This combination effectively captures spatial and temporal dependencies simultaneously within frames. With adversarial training, our method effectively reduces the unwanted artifacts in predictions, thereby ensuring the learning of more realistic and fluid human motions. The evaluation results indicate that AdvMT greatly enhances the accuracy of long-term predictions while also delivering robust short-term predictions.
@@ I Introduction
Human motion prediction is at the forefront of integrating disciplines such as artificial intelligence, robotics, and biomechanics, aiming to understand and forecast the complex and dynamic movements of humans. This area of research has broad applications, ranging from enhancing immersive human experience in virtual reality, to assistive robotics in healthcare and advanced manufacturing settings. However, the task of predicting human motion is challenging due to a multitude of influencing factors, including individual physical differences, psychological states, and the surrounding environment. These factors collectively complicate the development of realistic and reliable human motion prediction systems, a critical step for ensuring the smooth and harmonious integration of technology into human-centric applications.
@@ II-A Long-term human motion prediction
The initial phase of research in the area of human motion prediction primarily utilized RNNs with encoder-decoder models [1, 2, 3], but these models faced challenges in handling complex human motion dependencies. This led to a shift towards CNNs, which provided improved extraction of spatial joint connection information [5, 14, 6]. Later, the exploration of GCNs brought enhanced capabilities in anatomical relationship modeling [7, 8, 15]. However, much of the research has concentrated mainly on short-term predictions, leaving a notable gap in long-term prediction accuracy.
@@ III Adversarial Motion Transformer (AdvMT)
The overall architecture of the Adversarial Motion Transformer (AdvMT), a system comprised of two main branches, is illustrated in Fig. 2. The motion encoder branch interprets the input motion history and encodes the local and global human joint dependencies. The discriminator branch complements this by refining the predictions from the motion encoder branch to ensure the generation of realistic and consistent human motion.
@@ IV Experiments
Dataset: In our experiments, we conducted our model evaluation using the Human3.6M dataset, widely recognized as a benchmark in the field of human motion prediction. This extensive database contains over 3.6 million 3D poses, recorded with 7 actors performing 15 different types of actions. For training and evaluation, we downsampled the motion data to 25 frames per second. In alignment with the protocol outlined in [2], we used subjects S1, S6, S7, S8, S9, and S11 for training, and S5 is designated for testing.
Comparison with other methods: In our study, we focused on training our method for 3D joint position prediction. We compared its performance with other state-of-the-art methods trained for the same motion representation, as detailed in Tables I and II. The primary metric used for evaluation is the MPJPE in millimeters, aligning with the standards used by other methods [8, 19, 18, 20]. We trained the model using an input sequence of 2 seconds to generate predictions for the next 1 second. However, our motion encoder branch was trained auto-regressively, enabling it to predict motions extending beyond 1 second.
@@ V-A Architecture
An ablation study was conducted to assess the effectiveness of our proposed AdvMT architecture. We compared the performance of a full Transformer network against our modified architecture. Originally intended for sequence-to-sequence tasks, the full Transformer architecture was found less effective for human motion prediction compared to our adaptation, which solely utilizes the encoder layer. This suggests that the decoder layer might add unnecessary complexity, hindering the ability to capture human motion dynamics accurately.
Further ablation studies highlighted the critical role of the discriminator branch in enhancing the realism and temporal consistency of generated motion. The discriminator branch functions as a critic, ensuring the predicted motions are not only realistic but also align with real-world motion patterns. Its absence leads to the generation of unrealistic or inconsistent motion sequences, particularly in complex human motions where accuracy in joint angles and movements is crucial. By integrating the adversarial training, our model is compelled to generate more lifelike and consistent motion sequences.
@@ V-B Loss function
Regarding the loss function, we found that using only the bone length error as a loss term with the vanilla MPJPE loss led the model to predict zero-velocity motion for long-term predictions. As the model learns to minimize only the bone length error, leading it to generate static poses. To mitigate this, incorporating the discriminator loss term helped to penalize unrealistic human motion predictions and encouraged the motion encoder to generate more realistic and plausible human motion. We conducted an ablation study to investigate the effectiveness of our modified loss function, which includes a discriminator loss and a bone error loss in addition to the regular MPJPE loss. We compared the performance of our model with the modified loss function against a baseline model that only used the regular MPJPE loss. The results show that our modified loss function, which includes the MPJPE loss, bone length error, and a temporal continuity discriminator loss, is effective in improving the quality of human motion prediction.
TABLE III: The ablation study results on H3.6M dataset.